Resistance of a wire is y ohm.the wire is stretched to triple it's length,then the resistance becomes
1
Expert's answer
2021-09-16T14:19:02-0400

initial length = L

final length = 3L

initial resistance = y

"R_2 =n^2R_1 \\

R_2 = 3^2y \\

R_2 = 9y \; ohm"